OpenStudy (bha2908):

given- cos^4θ÷cos^2∅+sin^4θ÷sin^2∅=1 prove that cos^4∅÷cos^2θ+sin^4∅÷sin^2θ=1

jhonyy9 (jhonyy9):

cos^4 theta = cos^2 theta *cos^2 theta sin^4 theta = sin^2 theta *sin^2 theta rewrite these equation using above wrote and simplify and will get cos^2 theta +sin^2 theta = 1 what is true sure

OpenStudy (sshayer):

\[\theta~and~\phi ~are ~different~angles.\]